Write an equation of the line containing (9,-9) and parallel to the graph of 5x-9y=4.
The equation of the line is _=126

Answer by stanbon(75887)

Write an equation of the line containing (9,-9) and
parallel to the graph of 5x-9y=4
Given line in slope-intercept form:
y = (5/9)x - (4/9)
------
slope = 5/9
----
Form:
y = mx + b
Solve for "b":
-9 = (5/9)9 + b
b = -14
-----
Equation:
y = (5/9)x - 14
======================
Cheers,
Stan H.
===========================